OpenStudy (anonymous):

There are 20 students in a class. Determine how many different pairs you could make out of the class. Be sure to state if the problem involves a permutation or a combination and state if repetition is allowed. Show all of your work and explain your answer.

OpenStudy (anonymous):

there are lots of different ways to do this problem where did it come from? maybe we can pick a good strategy

OpenStudy (anonymous):

Its in my combination permutation lesson

OpenStudy (anonymous):

ok i would not imagine using any formula for this

OpenStudy (anonymous):

pick a student how many students can you pair her up with?

OpenStudy (anonymous):

Theres alot of ways.. have pairs of 2, 4, 5 & 10

OpenStudy (anonymous):

hold on a sec pick a student the question is "how may other students can you pair her up with?"

OpenStudy (anonymous):

is it clear that there are 19 other students you can pair her up with?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

Determine how many different pairs you could make out of the class with 20 students. It dont sday with just 1

OpenStudy (anonymous):

i think maybe you have misinterpreted the question, unless i have pairs mean 2 right? now "how many ways can you divide the class"

OpenStudy (anonymous):

yes i know, but we are not done just getting there

OpenStudy (anonymous):

with the first student, there are 19 ways then there are 18 students left pick one there are 17 other students to pair her up with now there are 16 left pick one there are 15 students to pair her up with etc

OpenStudy (anonymous):

by the counting principle, the number of ways to do all of this together is \[19\times 17\times 15\times 13\times 11\times 9\times 7\times 5\times 3\]

OpenStudy (anonymous):

654729075 is what i got

OpenStudy (anonymous):

Now what?